# Should You Use Your Large Language Model to Explore or Exploit?

### v2

@@ 2 LLMs as exploitation oracles
Our intuition for these findings is as follows: While LLMs are generally good with textual inputs, the reward data is numerical, so the LLM faces most/all of the same challenges as before (i.e.,  choosing a “solution strategy” and executing it correctly), exacerbated by non-linearity. Meanwhile, linear regression is known to often work fairly well even on non-linear data.
@@
Takeaways. While unmitigated performance drops as $d,K,T$ grow, increases in $T$ can largely be taken care of by our mitigations, whereas increases in $d$ and $K$ cannot. Indeed, taken together, our findings in this section suggest that while the current generation of LLMs are effective at exploiting in small-scale decision-making tasks, they struggle in larger settings.
@@ 3 LLMs as exploration oracles
Figure 6: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{task},K)$ (averaged over rounds and over runs), against $K$, the number of candidates. Each line corresponds to a prompting strategy or the $\mathtt{Random}$ baseline. The shaded regions represent a $95\%$ confidence interval.
@@
Our usage of LLMs for exploration is closely aligned with the idea of “discretization” in the study of dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011; Slivkins, 2014). Faced with a very large action space, it may be advantageous to focus on a much smaller set of candidate actions. A naively constructed discretization, e.g., uniform with respect to some known notion of distance in the action space, may have a poor tradeoff between size (#candidates) and quality (of the best candidate), placing all but a few candidates in low-performing regions of the action space. A crucial theme in this literature is “smarter” discretization algorithms which gradually “zoom in” on better-performing regions of the action space. The intuition is that if the better-performing arms comprise a low-dimensional region of a high-dimensional action space, covering this region should require much fewer samples for the same “discretization quality”. Our hope is that LLMs can suggest candidate actions from this region, if there is sufficient semantic meaning encoded in the bandit task.
We consider two types of exploration tasks: answering an open-ended “philosophical” question (Q/A task) and suggesting a title for an arXiv research paper based on its abstract (arXiv task). Particular workloads within each task type are called explore puzzles.
The Q/A task (resp. arXiv task) is constructed as follows. We define the “best arm” $a^{*}$ as a contrarian answer generated by another LLM (resp. the actual title of the research paper). The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between $a$ and $a^{*}$ in the embedding space. [^8] While cosine similarity ranges on [1,1][-1,1], it was usually strictly positive in our experiments. In the (very rare) cases where it was negative we defined the expected reward as zero. Here, we generate sentence embeddings using the Sentence-BERT embedding model Reimers (2019). [^9] Our results are largely unchanged under the Universal Sentence Encoder Cer (2018), see Appendix B. The realized reward in a given round is an independent Bernoulli sample with the mean defined above.
@@
Since action spaces are extremely large for these tasks, standard bandit algorithms (e.g.,  Thompson Sampling, UCB1, etc.) catastrophically fail, and so do bandit algorithms which randomly subsample the action space. If the embedding space, the distance notion therein, and the reward-distance relation were known to the algorithm/agent (assumptions we do not make), one could, in principle, apply the “smarter” discretization algorithms mentioned above. However, the regret of these algorithms scales exponentially in the dimension $d$ of the metric space making them impractical in high-dimensional action spaces such as ours. (Indeed, our embedding space has $d=384$.)
Like for our exploit puzzles, we highlight our results for GPT models in the main body and include the rest in Appendix B.5. Our findings are largely consistent across different LLMs.
Explore puzzle: open-ended questions. We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth. (E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, as defined above.
We evaluate Gpt-4o, Qwen-2.5, Gemma-3, and Mistral-7B as exploration oracles. We prompt them to suggest $K\in\{1,2,3,4,5,7,10\}$ candidate answers given a question. To grade the entire candidate set (not just the best answer), we run an off-the-shelf MAB algorithm for some large-but-realistic time horizon $T$ over these $K$ candidates. (We use the UCB1 algorithm (Auer et al., 2002) and $T=1000$.) We record the expected time-averaged reward $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, where $a_{t}$ is the arm selected in round $t$. We repeat this process (candidate selection and running UCB1) 10 times for a given task and $K$, and record the average $\mathtt{rew}$ over these runs, denoted by $\overline{\mathtt{rew}}=\overline{\mathtt{rew}}(\text{task},K)$.
Figure 7: Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right). We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run”, the $K$ suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
@@
We try several prompting strategies. We either ask the LLM to generate suggestions “all-at-once” with temperature 0 or “one-by-one” with temperature $1$ (we repeatedly show the LLM the list of candidate answers so far and ask it to generate one more). We also experiment with explicitly prompting the LLM to provide a diverse set of candidate answers (“with encouragement”). Thus, we have 4 prompting strategies: $\left\{\,\text{all-at-once, one-by-one}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement. To compare against not using an LLM as an exploration oracle, we also consider a baseline ($\mathtt{Random}$) in which the candidates are $K$ points selected independently and u.a.r. in the embedding space.
We visualize our findings in Figure 6. We plot $\overline{\mathtt{rew}}(\text{task},K)$ against $K$, for one particular task (left) and averaged across all tasks (right). [^10] See Appendix B for similar plots for the $9$ other tasks. Each line corresponds to a particular prompting strategy or the $\mathtt{Random}$ baseline. We find that all four prompting strategies perform similarly, with average reward between $0.5$ and $0.6$, and typically peak in performance around $K=3$ or $K=4$ suggestions. In contrast, the $\mathtt{Random}$ baseline catastrophically fails, with its average reward never exceeding $0.1$. We conclude that the LLM does succeed as an exploration oracle.
We observe that the LLM-generated suggestions pass the “eye test”, in the sense that we get reasonable, yet spiritually and semantically different answers for a given question. For example, given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our one-by-one prompt are as follows: (1) Facilitates communication, innovation, and efficiency, (2) Transforms daily life and shapes culture, (3) Drives connectivity and enhances productivity, (4) Facilitates control and surveillance, (5) Disrupts traditional relationships and norms.
Finally, we verify that the candidate suggestions are substantially different from one another. Figure 7 visualizes the spread of expected rewards within the candidate set. We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run” of the candidate selection, the suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
A more detailed comparison between the prompting strategies is not fruitful for these experiments, as sentence embeddings are known to be somewhat imprecise at small scales (see Section B.4).
@@
Figure 8: Algorithm’s average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates. Each line corresponds to a prompting strategy or the Category-Only baseline. The shaded regions represent a $95\%$ confidence interval. A single arXiv category (“General Relativity and Quantum Cosmology”, left), averages over 6 categories (right).
@@
Explore puzzle: arXiv abstracts and titles. We run similar experiments on a larger-scale dataset of paper titles and abstracts from arXiv.org. Using the arXiv API (arXiv.org, 2025), we collect $10$ abstract-title pairs from each of the $41$ different arXiv categories. To minimize the likelihood that these papers appear in Gpt-4o’s training corpus, we only use papers uploaded after June 2024 (these experiments were done in Sept’24). Each abstract-title pair yields a task, as discussed earlier.
We evaluate Gpt-4o as an exploration oracle for these tasks, much like for the Q/A tasks Given an abstract, we prompt Gpt-4o to generate $K$ alternative titles, which are then used to instantiate a bandit algorithm. We use the same algorithm (UCB1) and time horizon $T=1000$. We record the expected time-averaged reward, $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$, and compute the average over tasks within the same arXiv category, $\overline{\mathtt{rew}}(\text{category},K)$. We try “all-at-once” and “one-by-one” prompting. [^11] We do not use “encouragement” on these tasks, since it does not help much (if at all) for the Q/A tasks.
@@
We visualize our findings in Figure 8 , using the same conventions as Figure 6 and focusing on six arXiv categories. [^12] General relativity, computer vision, statistics theory, biomolecules, signal processing, general economics. To assess LLM’s ability to specialize to a task, we consider a stronger baseline, where the candidate arms are generated by Gpt-4o given only the category, not the abstract. For the other arXiv categories, we provide a table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K=1,2,5$ and both prompts (Appendix B). Like for the Q/A tasks, (1) both prompting strategies significantly outperform the baseline, and (2) their performance tends to improve modestly as $K$ increases.
@@ 4 Conclusions
Our work adds to (and provides partial explanations for) the growing literature on in-context learning for decision-making tasks. We find that LLMs are useful as exploration oracles that propose high-quality candidate actions in large, semantically meaningful action spaces. However, current LLMs are not that good at exploitation, particularly in larger or more complex tasks (we trace the “boundary of tractability” in terms of task size/complexity). While we suggest several helpful mitigations, they consistently underperform relative to a simple linear regression, even in inherently non-linear tasks.
Future work. LLMs trained to use tools like a calculator may be better at exploitation. However, it is unclear how much this would help in more complex scenarios, e.g.,  CB tasks with text-based contexts and actions, and which mitigations/prompting techniques would be needed. Second, while ”zooming” bandit algorithms do not work for rich text-based action spaces (discussed in Section 3), LLM-based exploration oracles may potentially help. The hope is to “zoom in” entirely in the space of “potentially relevant” actions (as determined by the LLM), rather than in the space of all actions.
Appendix A Appendix for Section 2: LLMs as exploitation oracles
A.1 Prompts
@@
A.2 Additional MAB figures
See Figure 9Figure 10Figure 11 for additional results in our MAB exploit puzzle.
@@ A.2 Additional MAB figures
A.3 Additional details for text-based CB puzzles
Each context contains a time of day (belonging to {morning, afternoon, evening, night}), an animal ({bear, dog, cat, None}), a tool ({key, letter opener, hammer, None}), a food item ({cake, apple, nut, None}), and a button with a particular color ({red, orange, yellow, green}). The actions in each round are “pet animal”, “leave room”, “use tool”, “eat food”, and “press button”.
We experimented with two reward functions: an “easy” reward function, where the expected rewards for each action are as follows:
  • The expected reward of petting the animal is 0.01 if the animal is a bear, 0.7 if the animal is a dog, and 0.4 if the animal is a cat. Otherwise, the expected reward if 0.5.
  • The expected reward for leaving the room is always 0.5.
  • The expected reward for using the tool is 0.75 if it is a key, 0.6 if it is a letter opener, 0.45 if it is a hammer, and 0.2 otherwise.
  • The expected reward for eating food is 0.8 if it is cake, 0.6 if it is an apple, 0.2 if it is a nut, and 0.3 otherwise.
  • The expected reward for pressing the button is 0.89 if it is green, 0.62 if it is yellow, 0.39 if it is orange, and 0.27 if it is red.

### v3

@@ 2 LLMs as exploitation oracles
Our intuition for these findings is as follows: While LLMs are generally good with textual inputs, the reward data is numerical, so the LLM faces same challenges as before (choosing a “solution strategy” and executing it correctly), exacerbated by non-linearity. Meanwhile, linear regression is known to often work fairly well even on non-linear data.
Takeaways. Frontier models can now solve fairly complicated MAB exploit puzzles. In contrast, Gpt-4 and Gpt-3.5 struggle, partially explaining their failure on end-to-end MAB tasks in Krishnamurthy et al. (2024); Nie et al. (2024).
@@
In numerical CB puzzles, unmitigated performance drops as $d,K,T$ grow, but tool use can help and increases in $T$ can largely be taken care of by our in-context mitigations. In contrast, our mitigations were largely unhelpful in a non-linear, text-based CB puzzle. Taken together, our findings suggest that while current LLMs are effective at exploiting in small-scale, numerical decision-making tasks, they still struggle in more complicated settings.
Figure 5: Performance of the Gpt-5.2 (left) and Gpt-4o on a text-based CB puzzle. In this more challenging non-linear setting, some mitigations help, but all are outperformed by linear regression.
@@ 3 LLMs as exploration oracles
This usage of LLMs is closely aligned with “discretization” in dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011; Slivkins, 2014). Given a very large action space, it may help to focus on a much smaller set of candidate arms. A naively constructed discretization, e.g., uniform w.r.t. some known distance notion on the arms, may have a poor tradeoff between size (#candidates) and quality (of the best candidate), placing all but a few candidates in low-performing regions of the action space. A crucial theme in this literature is “smarter” discretization algorithms which gradually “zoom in” on better-performing arms. If the latter comprise a low-dimensional region of the action space, covering this region should require much fewer samples for the same “discretization quality”. Our hope is that LLMs can suggest candidate actions from this region, if there is sufficient semantic meaning encoded in the bandit task.
@@
Non-LLM approaches appear hopeless for our tasks. Standard bandit algorithms (Thompson Sampling, UCB1, etc.) fail for a very large #arms, and so do bandit algorithms which randomly subsample the arms. Another approach embeds the arms’ text labels in $\mathbb{R}^{d}$. If a suitable reward-distance relation were known to the algorithm/agent (assumption we do not make), one could apply the “smarter” discretization algorithms mentioned above. However, their regret scales exponentially in $d$, making them impractical for state-of-art text embeddings (e.g., one we use has $d=384$).
Our tasks. We consider stylized exploration tasks of the following form. In each task, we define an MAB instance with Bernoulli rewards and a very large #arms, each labeled with a text string. The mean rewards $\mu(\cdot)$ and the meaning of the arms’ labels are based on real-world data. We call an LLM once to generate a small representative subset of $K$ arms. To grade the entire set (not just the best arm therein), we run a standard bandit algorithm (UCB1; Auer et al. (2002)) over these $K$ arms, for $T=1000$ rounds. We record the algorithm’s average cumulative reward, “denoised” as $\mathtt{rew}:=\tfrac{1}{T}\,\sum_{t\in[T]}\mu(a_{t})$.
We consider three task types: (1) recommend a movie, based on the MovieLens 100K dataset Harper and Konstan (2015), a standard benchmark for evaluating CB (MovieLens task), (2) suggest a title for an arXiv research paper given its abstract (arXiv task), and (3) answer an open-ended “philosophical” question (Q/A task). Particular workloads within each task type are called explore puzzles. Our findings are similar across all task types and all LLMs.
The MovieLens 100K dataset contains ratings from 943 users across 1,682 movies. In our tasks, arms correspond to the movies titles. Each task corresponds to a particular user, with mean rewards given by this user’s ratings. Since every user did not rate every movie, we used a standard algorithm to complete the data matrix (the soft-impute low-rank matrix completion method from the fancyimpute Python package). We then normalize this matrix so that all values lie $[0,1]$.
The arXiv task (resp. Q/A task) is constructed as follows. Each task corresponds to an abstract-title pair from arXiv (resp., a question-answer pair). The arms are all possible titles (resp., answers). The “best arm” $a^{*}$ is the actual title (resp. a contrarian answer generated by another LLM). The expected reward $\mu(a)$ of an arm $a$ is the cosine similarity between $a$ and $a^{*}$ in the embedding generated by the Sentence-BERT embedding model Reimers (2019).  [^8] While cosine similarity ranges on [1,1][-1,1], it was usually strictly positive in our experiments. In the (very rare) cases where it was negative we defined the expected reward as zero. [^9] Our results are largely unchanged under the Universal Sentence Encoder Cer (2018), see Section C.4.
@@
As a robustness check, we try several prompting strategies (detailed in Appendix C.1). We either ask the LLM to generate suggestions “all-at-once” with temperature 0 or “1-by-1” with temperature $1$ (we repeatedly show the LLM the candidate arms so far and ask it to generate one more). We also try to explicitly encourage the LLM to provide a diverse candidate set. Thus, we have 4 prompting strategies: $\left\{\,\text{all-at-once, 1-by-1}\,\right\}\times\left\{\,\text{with, without}\,\right\}$ encouragement. However, all 4 strategies performed similarly. [^10] We only tried the 4 strategies for the arXiv and Q/A tasks.
@@
Figure 6: Average expected reward $\overline{\mathtt{rew}}(\text{category},K)$ (averaged over rounds and tasks), against $K$, the number of candidates. Each line corresponds to a prompting strategy or the Category-Only baseline. The shaded regions represent a $95\%$ confidence interval. A single arXiv category (“General Relativity and Quantum Cosmology”, left), averages over 6 categories (right).
Explore puzzle: MovieLens. We run a MovieLens task for each of the $943$ users, with $K\in\{10,19\}$. We re-run each task 5 times to denoise. For a given $K$, we record $\overline{\mathtt{rew}}(K)$: reward $\mathtt{rew}$, defined above, averaged over all users and runs. For baselines, we compare against running UCB1 for $T$ rounds over (1) a random subset of $K$ movies and (2) a random movie from each of the $19$ movie genres in the dataset. (Different random draws for each user.)
Our results for Mistral-7B and Qwen-2.5 are in Table 1 Both LLMs significantly outperform all baselines for both values of $K$. We also tried running the experiment using Llama-3-8B and Gemma-3, but the list of all movies was too long to fit in the context window.
Table 1: Average Reward for Qwen-2.5, Mistral-7B, and Baselines for $K\in\{10,19\}$ in the MovieLens Task.
$\overline{\mathtt{rew}}$ $K=10$ $K=19$
Qwen-2.5 $671.88$ $\mathbf{678.25}$
Mistral-7B $\mathbf{735.38}$ $675.16$
Random $570.39$ $599.73$
Genre-based $617.27$
The LLM has implicit knowledge about each movie which it may leverage when selecting the discretization. To mitigate overfitting to “popular” movies, we prompted the LLM to not take movie popularity into account. However, this change did not significantly affect the rewards.
Table 2: Performance using all-at-once prompting on select arXiv tasks. Results are averaged over $10$ abstract-title pairs. The full table (Table 3) may be found in the Appendix.
$K=1$ $K=2$ $K=5$
gr-qc 0.63 0.64 0.68
hep-ex 0.81 0.81 0.83
hep-th 0.65 0.71 0.73
math-ph 0.64 0.73 0.74
nucl-ex 0.73 0.79 0.75
@@
Explore puzzle: arXiv. Using the arXiv API (arXiv.org, 2025), we collected $10$ abstract-title pairs from each of the $41$ arXiv categories. [^11] To minimize the likelihood that these papers appear in the training corpus, we only use papers uploaded after June’24 (these experiments were done in Sept’24). Each abstract-title pair yields a task, with $K=1,2,5$ suggestions. Each task is re-run 5 times to denoise. Given $K$, we record $\overline{\mathtt{rew}}(\text{category},K)$: reward $\mathtt{rew}$, as defined above, averaged over all tasks within the same arXiv category and all runs.
@@
We evaluate Gpt-4o as an exploration oracle for these tasks. We visualize our findings in Figure 6 , focusing on six arXiv categories. To assess the LLM’s ability to specialize to a task, we consider a baseline where the candidate arms are generated by Gpt-4o given only the category, not the abstract. For the other arXiv categories, we provide a table with $\overline{\mathtt{rew}}(\text{category},K)$ for $K=1,2,5$ (Table 2, Appendix C).
Explore puzzle: Q/A task. We used Gpt-4 to generate a dataset of $10$ open-ended questions with many reasonable answers, along with an intentionally contrarian answer for each question to serve as the ground truth. (E.g.,  “What does it mean to live a fulfilling life?” “Fulfillment comes from embracing discomfort.”) Each question-answer pair yields a task, with $K\in\{1,2,3,4,5,7,10\}$ candidate answers. For a given $K$, we run each task 10 times and record $\overline{\mathtt{rew}}(\text{task},K)$: reward $\mathtt{rew}$ averaged over all runs.
We evaluate Gpt-4o, Qwen-2.5, Gemma-3, and Mistral-7B as exploration oracles. We also consider a non-LLM baseline ($\mathtt{Random}$), which selects $K$ arms independently and u.a.r. in the embedding space (leveraging the AI-based embedding but suffering from its high dimension).
The LLMs perform similarly, with $\overline{\mathtt{rew}}\approx$0.5-0.6. and typically work best for KK=3,4. see Appendix A. In contrast, $\mathtt{Random}$ catastrophically fails, with $\overline{\mathtt{rew}}<0.1$. Moreover, we observe that LLM-generated answers pass the “eye test” (they are reasonable, yet spiritually and semantically different), and yield substantially different expected rewards. We conclude that the LLM does succeed as an exploration oracle for these tasks.
@@ 4 Conclusions
Our work adds to (and provides partial explanations for) the growing literature on in-context learning for decision-making tasks. We find that LLMs are useful as exploration oracles that propose high-quality candidate actions in large, semantically meaningful action spaces. However, current LLMs are not that good at exploitation, particularly in larger or more complex tasks, although reasoning models (specifically, the Gpt-5 family of models) show the most promise. While we experiment with several mitigations including tool use and in-context summaries, they consistently underperform relative to a simple linear regression baseline, even in inherently non-linear tasks.
We provide a systematic evaluation of LLMs’ capabilities as exploration oracles, in isolation from other components of a decision-making agent. We find robust success across several different task types, LLMs, and prompting strategies. In contrast, non-LLM baselines are often quite weak, absent known problem structure to be leveraged by an algorithm.
LLMs’ continued difficulties at exploitation tasks should be contrasted with the rapid progress in LLMs and the extreme optimism associated with this progress. Ever since the Gpt-3.5 release in 2022 (and especially since Gpt-4 frenzy of early 2023), it was commonly believed that “generalist” frontier LLMs will soon master many/most relatively simple tasks: if not yet, then more likely in several months than in several years. However, our results show that these difficulties persist longer than anticipated.
Future work. LLM-based exploration oracles may potentially help with “smart discretization” for rich text-based action spaces. Non-LLM bandit algorithms tend to fail in such scenarios, as discussed in Section 3. The hope is to “zoom in” entirely in the space of “potentially relevant” actions, rather than in the space of all actions.
Second, while we experiment with tool use for exploitation tasks, we find that it does not improve performance across the board. One could investigate various approaches for tool use in these tasks, and whether/when they help.
Acknowledgements
We would like to thank Dylan Foster and Akshay Krishnamurthy for helpful conversations at early stages of the project.
Appendix A Main Plots for the Q/A explore task
Here we provide the key plots for the Q/A explore task, as defined in Section 3. As stated there, all LLMs we considered perform similarly, with $\overline{\mathtt{rew}}\approx$0.5-0.6. and typically work best for KK=3,4, whereas the $\mathtt{Random}$ baseline catastrophically fails, with $\overline{\mathtt{rew}}<0.1$. Recall that $\overline{\mathtt{rew}}(\text{task},K)$ is the average reward $\mathtt{rew}$ for a given task with a given $K$ (as defined in Section 3), averaged over all runs of this task.
We visualize our findings in Figure 7. We plot $\overline{\mathtt{rew}}(\text{task},K)$ against $K$, for one particular task (left) and averaged across all tasks (right). Similar plots for the $9$ other tasks can be found in Section C.6.
Figure 7: Average expected reward $\overline{\mathtt{rew}}(\text{task},K)$ (averaged over rounds and over runs), against $K$, #candidates. Each line corresponds to a prompting strategy or the $\mathtt{Random}$ baseline. The shaded regions represent a $95\%$ confidence interval.
The LLM-generated answers pass the “eye test”: they are reasonable, yet spiritually and semantically different. E.g., given the question “What is the role of technology in society”, the first $K=5$ suggestions generated by our 1-by-1 prompt are: (1) Facilitates communication, innovation, and efficiency, (2) Transforms daily life and shapes culture, (3) Drives connectivity and enhances productivity, (4) Facilitates control and surveillance, (5) Disrupts traditional relationships and norms. Finally, we verify that the candidate suggestions are substantially different from one another in terms of their rewards, see Figure 8.
Figure 8: Spread of expected rewards within the candidate set. Arm histogram for one particular question (“What is the meaning of freedom?”, left) and averaged over all questions (right). We consider $K=10$ suggestions generated by our one-by-one prompt without encouragement. For a given “run”, the $K$ suggestions are ranked by the expected reward, and then the $i$-th expected reward is averaged over all runs, for each $i\in[K]$.
Appendix B Details for Section 2: LLMs as exploitation oracles
B.1 Prompts
@@
B.2 Additional Results on MAB exploit puzzle
@@
B.3 Additional details for text-based CB puzzles
Each context contains a time of day (belonging to {morning, afternoon, evening, night}), an animal ({bear, dog, cat, None}), a tool ({key, letter opener, hammer, None}), a food item ({cake, apple, nut, None}), and a button with a particular color ({red, orange, yellow, green}). The actions are “pet animal”, “leave room”, “use tool”, “eat food”, and “press button”.
We considered two reward functions. One was “easy”, with expected rewards (ExpRev) as follows:
  • ExpRev for petting the animal is 0.01 if the animal is a bear, 0.7 if the animal is a dog, and 0.4 if the animal is a cat. Otherwise, ExpRev if 0.5.
  • ExpRev for leaving the room is always 0.5.
  • ExpRev for using the tool is 0.75 if it is a key, 0.6 if it is a letter opener, 0.45 if it is a hammer, and 0.2 otherwise.
  • ExpRev for eating food is 0.8 if it is cake, 0.6 if it is an apple, 0.2 if it is a nut, and 0.3 otherwise.
  • ExpRev for pressing the button is 0.89 if it is green, 0.62 if it is yellow, 0.39 if it is orange, and 0.27 if it is red.